OpenStudy (anonymous):

juanita and lauran are painting a circular mural on one wall of the community center. the area of the mural is 88 square meters. what is the radius? use pi=22/7 the answer wud be 11 right

OpenStudy (luigi0210):

@Jhannybean

OpenStudy (anonymous):

.... ummm

OpenStudy (jhannybean):

Do you know the formula for the area of a circle?

OpenStudy (anonymous):

pi equals radius times c

OpenStudy (jhannybean):

O_o

OpenStudy (anonymous):

haha idk :( srry

OpenStudy (jhannybean):

xD \(\large A = \pi r^2\) For a circle Givens \(A = 88 \ m^2\) and \( \pi = \cfrac{22}{7} \) Let's change our ewuation around to solve for the radius :D \[\large r^2 = \cfrac{A}{\pi}\]\[\large r = \pm \sqrt{\cfrac{A}{r}}\]\[\large r = \pm \sqrt{ \cfrac{88}{\cfrac{22}{7}}} = \pm \sqrt{28} \]

OpenStudy (anonymous):

o i see now

OpenStudy (jhannybean):

omg. the second line from the bottom should have said \[\large r = \pm \sqrt{\cfrac{A}{\pi}} \]

OpenStudy (jhannybean):

That was a typo...

OpenStudy (anonymous):

oh hahah i c it nowe soo my answer isnt 11

OpenStudy (jhannybean):

Yeah :D

OpenStudy (anonymous):

can u help meh with one more??

OpenStudy (jhannybean):

Not atm :( I've got to head off OS for a little bit. I'll be afk D:

OpenStudy (anonymous):

o okaii bye :)

OpenStudy (jhannybean):

Bye bye! Good luck :)

OpenStudy (anonymous):

you too